seven movies (A, B, C, D, E, F, and G) are being scheduled for showing. The order of the showing is determined by random relecton. find the probability taht film C will be shown first, film A next-to-last, and film E last.
Answer by stanbon(75887)

seven movies (A, B, C, D, E, F, and G) are being scheduled for showing. The order of the showing is determined by random selecton. find the probability that film C will be shown first, film A next-to-last, and film E last.
--------------------
# of ways to select 1st film: 1
# of ways to select 6th film: 1
# of ways to select 2,3,4,5, and 7th film: 5! = 120
--------------------------
# of ways to get the pattern you want: 1*1*120 = 120 ways
--------------
# of ways to make the selections without restriction: 7! = 5040
-----------------------------------------------
Probability = 5!/7! = 1/(6*7) = 1/42
=================================
Cheers,
Stan H.